Sony would like to test the hypothesis that the average age of a PlayStation user is different from the average age of an Xbox user. A random sample of 36 PlayStation users had an average age of 34.2 years while a random sample of 30 Xbox users had an average age of 32.7 years. Assume that the population standard deviation for the age of PlayStation and Xbox users is 3.9 and 4.0​ years, respectively. Sony would like to set alpha equals 0.10.

What is the​ p-value for this hypothesis​ test? Round to four decimal places as needed.

A. 0.1260

B. 0.0319

C. 0.0643

D. 0.1722
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: μ1 = μ2
Alternative Hypothesis, Ha: μ1 ≠ μ2


Pooled Variance
sp = sqrt(s1^2/n1 + s2^2/n2)
sp = sqrt(15.21/36 + 16/30)
sp = 0.9777

Test statistic,
z = (x1bar - x2bar)/sp
z = (34.2 - 32.7)/0.9777
z = 1.53

P-value Approach
P-value = 0.1260


A. 0.1260